# Every SAM Drop Counts: Embracing Semantic Priors for Multi-Modality Image Fusion and Beyond

### v1

@@ 4 Experiments
Table 1: Quantitative semantic segmentation IoU results of different methods on the FMB dataset.
Figure 6: Visual comparison of segmentation results of different methods on the FMB dataset.
@@ 4.1 Settings and Details
Five representative datasets, namely TNO [34], RoadScene [42], MFNet [4], FMB [12], and M3FD [10], are utilized for both training and evaluation. For fusion-based segmentation evaluations, SegFormer [41] (B3 variant) is adopted as the backbone, and the models are trained for 100 epochs. The training and testing splits are strictly in accordance with the official dataset guidelines. The Adam optimizer is used to optimize both networks during the training period, with learning rates of $5\times 10^{-5}$ for both networks. Prior to distillation, the models undergo a pre-training phase, followed by 5 epochs of distillation. The batch size is set to 4, and the images are randomly cropped and resized to $192\times 160$ during training. The entire framework is implemented in PyTorch and executed on two NVIDIA GeForce RTX 4090 GPUs.
4.2 Results of IVIF
@@
Visualization Comparisons. Figure. 4 presents a visual comparison of various methods. Overall, the proposed method has two main advantages. First, it effectively preserves the multi-modality information of the original images. In the TNO surveillance scene, the vegetation details from the visible image and the smoke from the chimney in the infrared image are both well retained (the image set in the top-left corner). In RoadScene, our method also achieves optimal leaf recovery. On the other hand, our approach demonstrates strong robustness to interference, as it accurately reconstructs the reflective pedestrian crossing lines on the ground and the distant building outlines in the dense fog at night (the green box in the second row). The integration of SAM enhances our method, giving it the ability to surpass other methods and achieve superior results.
Quantitative Results. We also compare our numerical results with other competitors in Figure. 5, based on 100 image pairs randomly selected from the FMB and M3FD datasets. We employ four fusion evaluation metrics, including the Entropy (EN) [31], Standard Deviation (SD) [1], Sum of Correlation Differences (SCD) [24], and Multi-Scale Structural Similarity (MS-SSIM) [39]. Additionally, we introduce four No-Reference Image Quality Assessments (NR-IQA) related to image quality: BRISQUE [27], NIQE [28], MUSIQ [7], and PaQ-2-PiQ [43].
@@
Method Venue Car Per. Bike Curve C.S. C.C. Bump mIoU
Visible - 83.3 54.7 59.6 7.46 31.0 37.4 30.8 44.6
Infrared - 79.0 63.8 51.4 7.10 25.2 34.1 41.8 44.4
DDFM ICCV’23 86.7 62.2 63.4 30.8 35.5 45.7 41.1 51.7
U2Fusion TPAMI’22 87.3 65.9 60.8 33.4 36.0 46.1 40.9 52.0
TarDAL CVPR’22 86.4 67.1 62.4 28.4 30.5 45.0 41.5 51.0
SegMiF ICCV’23 85.9 66.3 64.4 34.0 34.7 43.0 48.0 52.4
FILM ICML’24 86.6 66.4 61.9 36.3 32.0 46.7 46.7 52.7
SHIP CVPR’24 86.7 63.9 59.7 33.5 35.2 44.4 42.4 51.5
MRFS CVPR’24 86.1 61.5 59.7 30.6 35.2 44.9 38.4 50.4
EMMA CVPR’24 86.1 65.6 63.8 25.4 29.1 44.9 41.3 51.7
TIMFusion TPAMI’24 86.4 58.7 60.3 29.8 33.6 44.1 38.3 49.8
SAGE Proposed 87.9 67.5 64.3 41.2 34.1 46.8 46.7 54.0
Table 2: Quantitative semantic segmentation results of different methods on the MFNet dataset.
Figure 7: Visual comparison of segmentation results of different methods on the MFNet dataset.
4.3 Results of IVIF based Segmentation
Visualization Comparisons. We present the segmentation visualization results on the challenging new FMB dataset in Figure. 6. In the daytime intersection scene, thanks to the powerful semantic priors provided by SAM, our method is the only one to completely distinguish between the truck and the bus. In the nighttime road scene, we successfully segment the sidewalk, achieving optimal performance. Additionally, we showcase the segmentation results on the MFNet dataset in Figure. 7. Leveraging the advanced semantic information learned by SAM, our method successfully segments distant small pedestrian targets during the daytime and near-invisible lane curve at night.
Quantitative Results. Table. 1 presents a qualitative comparison of IoU results for two segmentation frameworks on the FMB dataset. The first framework involves re-training images generated by various fusion methods, while the second uses the output from the open-vocabulary segmentation network with prompt words, without requiring retraining. In traditional comparisons, our method achieves a 3.0 mIoU improvement over the second-best approach, demonstrating its competitive edge across multiple categories. Additionally, in the segmentation network that requires no training, our method also performs favorably, benefiting from the efficient adaptability provided by SAM’s semantic information. Table. 2 further showcases the qualitative results on the MFNet dataset, which has lower resolution and fewer labels. Although our method shows some competitive gaps in background categories (e.g., bicycles, colored cones), it achieves significant overall superiority in key perception categories (e.g., Car and Person) and mean performance. Our method seamlessly integrates SAM with the fusion network, achieving a breakthrough in performance for efficient segmentation task.
Variants mIoU Metric
@@ II. Design of SPA
III. Distillation Triplet Losses
@@
(d) Online Dist. 50.0 / 50.4 6.221 / 32.65 / 1.581 / 0.987
The proposed method 54.0 / 61.2 6.872 / 40.77 / 1.604 / 1.117
@@ 4.4 Ablation Studies
Impact of SAM. The Segment Anything Module serves as a critical cornerstone in our method. To explore its impact, we derive three variants by isolating SAM from the core framework. Specifically, in variant (a), we replace the semantic patches with randomly cropped source image patches, thereby removing SAM from the main network $\mathcal{M}$. In variant (b), we assist the generation of semantic patches with segmentation labels. Additionally, we replace $\mathcal{S}^{\text{O}}$ with a conventional segmentation network [41]. Variant (a) enhances the source image information, leading to the optimal similarity metrics, SCD and MS-SSIM, as shown in Table. 3. Our method fully leverages the semantic priors provided by SAM, resulting in the best foggy building contours, as illustrated in the first row of Figure. 8.
@@
Study on SPA module. The proposed Semantic Persistent Attention module plays a key role in integrating the semantic priors from SAM, effectively guiding the retention and enhancement of inherent modality features. To investigate its impact, we set up three variants: (a) w/o latent representation (Z), (b) w/o key-value pairs in the Persistent Repository (PR), and (c) w/o PR. The channel loss due to these modifications has been corrected. A visual comparison of these ablation variants is shown in Figure. 9. It is evident that PR in the SPA module plays an essential role in maintaining critical information. The removal of any component leads to significant information loss. For instance, in (b), while the regions of interest are highlighted, the absence of source image information causes blurriness. Similarly, the absence of PR in the SPA module completely prevents the capture of beneficial semantic information, causing the network to lose focus on key areas and resulting in a low-contrast fusion results in (c).
Discussion on distillation scheme. Furthermore, we discuss the distillation scheme. We first conduct ablation studies on each component of the triple loss function, yielding variants (a)-(c) in Table. 3 III. We also replace the dual-layer optimization distillation with a traditional online distillation approach, forming variant (d). A visual comparison is shown in Figure. 10. Notably, traditional online distillation results are significantly inferior to our method, both in visual quality and the gradient map from variant (b). This confirms the effectiveness of our distillation method when incorporating semantic information.
@@ 4.5 Computational Efficiency Analysis
In the M3FD benchmark, we compare our method with 9 other SOTA methods in terms of time, FLOPS, and parameters. As shown in Table. 4, our method demonstrates significant advantages across all aspects, particularly in time and FLOPS. Specifically, due to our distillation scheme, our sub-network is able to adjust flexibly while maintaining high efficiency, resulting in reduced FLOPS. Compared to other methods, our approach achieves a processing time of 10.47 ms and FLOPS of 52.06 G, outperforming most of the existing methods, with a parameter count of only 0.136M, showcasing computational efficiency.

### v2

@@ 4 Experiments
Table 1: Quantitative semantic segmentation results of SOTA approaches on the FMB dataset.
Figure 6: Qualitative segmentation results of SOTA approaches on the FMB dataset.
@@ 4.1 Settings and Details
Five representative datasets, namely TNO [27], RoadScene [32], MFNet [2], FMB [13], and M3FD [11], are utilized for both training and evaluation. For segmentation, SegFormer [31] (B3 variant) is adopted as the backbone, and the models are trained for 100 epochs. The training and testing splits are strictly in accordance with the official dataset guidelines. Adam is adopted for training, with initial learning rates of $5\times 10^{-4}$ and $2\times 10^{-3}$ for the main and sub-networks, respectively. Cosine annealing decays both to $1\times 10^{-5}$. Prior to distillation, the models undergo a pre-training phase, followed by 5 epochs of distillation. The batch size is set to 4, and the images are randomly cropped and resized to $192\times 256$ during training. The entire framework is implemented in PyTorch and executed on two NVIDIA GeForce RTX 4090 GPUs.
4.2 Results of Multi-modality Image Fusion
@@
Qualitative Comparisons. Figure 4 presents a visual comparison of various methods. Overall, the proposed method has two main advantages. First, it effectively preserves the multi-modality information of the original images. In the TNO surveillance scene, the vegetation details from the visible image and the smoke from the chimney in the infrared image are both well retained (the image set in the top-left corner). In RoadScene, our method also achieves optimal leaf recovery. On the other hand, our approach demonstrates strong robustness to interference, as it accurately reconstructs the reflective pedestrian crossing lines on the ground and the distant building outlines in the dense fog at night (the green box in the second row). The integration of SAM enhances our method, giving it the ability to surpass other methods and achieve superior results.
Quantitative Comparisons. We also compare our numerical results with other competitors in Figure 5, based on 100 image pairs randomly selected from the FMB and M3FD datasets. We employ four fusion evaluation metrics [21], including the Entropy (EN), Standard Deviation (SD), Sum of Correlation Differences (SCD), and Multi-Scale Structural Similarity (MS-SSIM). Additionally, we introduce four no-reference image quality assessments [34] related to image quality: BRISQUE, NIQE, MUSIQ, and PaQ-2-PiQ.
@@
Method Car Per. Bike Curve C.S. C.C. Bump mIoU
Visible 83.3 54.7 59.6 7.46 31.0 37.4 30.8 44.6
Infrared 79.0 63.8 51.4 7.10 25.2 34.1 41.8 44.4
DDFM 86.7 62.2 63.4 30.8 35.5 45.7 41.1 51.7
U2Fusion 87.3 65.9 60.8 33.4 36.0 46.1 40.9 52.0
TarDAL 86.4 67.1 62.4 28.4 30.5 45.0 41.5 51.0
SegMiF 85.9 66.3 64.4 34.0 34.7 43.0 48.0 52.4
FILM 86.6 66.4 61.9 36.3 32.0 46.7 46.7 52.7
SHIP 86.7 63.9 59.7 33.5 35.2 44.4 42.4 51.5
MRFS 86.1 61.5 59.7 30.6 35.2 44.9 38.4 50.4
EMMA 86.1 65.6 63.8 25.4 29.1 44.9 41.3 51.7
TIMFusion 86.4 58.7 60.3 29.8 33.6 44.1 38.3 49.8
SAGE 87.9 67.5 64.3 41.2 34.1 46.8 46.7 54.0
Table 2: Quantitative semantic segmentation results of SOTA approaches on the MFNet dataset.
Figure 7: Qualitative segmentation results of SOTA approaches on the MFNet dataset.
4.3 Results of Multi-modality Segmentation
Qualitative Comparisons. We present the segmentation visualization results on the challenging new FMB dataset in Figure 6. In the daytime intersection scene, thanks to the powerful semantic priors provided by SAM, our method is the only one to completely distinguish between the truck and the bus. In the nighttime road scene, we successfully segment the sidewalk, achieving optimal performance. Additionally, we showcase the segmentation results on the MFNet dataset in Figure 7. Leveraging the advanced semantic information learned by SAM, our method successfully segments distant small pedestrian targets during the daytime and near-invisible lane curve at night.
Quantitative Comparisons. Table 1 presents a qualitative comparison of IoU results for two segmentation frameworks on the FMB dataset. The first framework involves re-training images generated by various fusion methods, while the second uses the output from an open-vocabulary segmentation network with prompt words, without retraining. In traditional comparisons, our method achieves a 3.0 mIoU improvement over the second-best approach, demonstrating its competitive edge across categories. Additionally, in the segmentation network that requires no training, our method also performs favorably, benefiting from the adaptability provided by SAM’s semantic information. Table 2 further showcases the qualitative results on the MFNet dataset, which has lower resolution and fewer labels. Although our method shows some gaps in background categories, it achieves overall superiority in key perception categories and mean performance. Our method integrates SAM with the fusion network, achieving a breakthrough in performance for efficient segmentation.
Variants mIoU Fusion Metrics
@@ II. Study on the SPA Module
III. Discussion on the Distillation Scheme
@@
(d) Offline Dist. 50.0 / 50.4 6.221 / 32.65 / 1.581 / 0.987
SAGE 54.0 / 61.2 6.872 / 40.77 / 1.604 / 1.117
@@ 4.4 Ablation Studies
The Impact of SAM. The Segment Anything Module serves as a critical cornerstone in our method. To explore its impact, we derive three variants by isolating SAM from the core framework. Specifically, in variant (a), we replace the semantic patches with randomly cropped source image patches, thereby removing SAM from the main network $\mathcal{M}$. In variant (b), we assist the generation of semantic patches with segmentation labels. Additionally, we replace $\mathcal{S}_{\text{O}}$ with a conventional segmentation network [31]. Variant (a) enhances the source image information, leading to the optimal similarity metrics, SCD and MS-SSIM, as shown in Table 3. Our method fully leverages the semantic priors provided by SAM, resulting in the best foggy building contours, as illustrated in the first row of Figure 8.
@@
Study on the SPA Module. The proposed Semantic Persistent Attention module plays a key role in integrating the semantic priors from SAM, effectively guiding the retention and enhancement of inherent modality features. To investigate its impact, we set up three variants: (a) w/o latent representation Z, (b) w/o key-value pairs in the Persistent Repository, and (c) w/o PR. The channel loss due to these modifications has been corrected. A visual comparison of these ablation variants is shown in Figure 9. It is evident that PR in the SPA module plays an essential role in maintaining critical information. The removal of any component leads to significant information loss. For instance, in (b), while the regions of interest are highlighted, the absence of source image information causes blurriness. Similarly, the absence of PR in the SPA module completely prevents the capture of beneficial semantic information, causing the network to lose focus on key areas and resulting in a low-contrast fusion results in (c).
@@
Discussion on the Distillation Scheme. Furthermore, we discuss the distillation scheme. First, we conduct ablation studies on each component of the triple loss function, leading to variants (a)-(c) as shown in Table 3 III. Additionally, we replace the bi-level optimization distillation method with the offline distillation approach, forming variant (d). A visual comparison of these variants is shown in Figure 10. Notably, the results of the traditional online distillation are significantly inferior to our method, both in terms of visual quality and the gradient map generated by variant (b). This confirms the effectiveness of the proposed distillation method when incorporating semantic information.
@@ 4.5 Computational Efficiency Analysis
In the M3FD benchmark, we compare our method with 9 other SOTA methods in terms of time, FLOPS, and parameters. As shown in Table 4, our method demonstrates significant advantages across all aspects, particularly in time and FLOPs. Specifically, due to our distillation scheme, our sub-network is able to adjust flexibly while maintaining high efficiency, resulting in reduced FLOPs. Compared to other methods, our approach achieves a processing time of 10.47 ms and FLOPs of 52.06 G, outperforming most of the existing methods, with a parameter count of only 0.136M, showcasing computational efficiency.
@@ 5 Conclusion
Acknowledgments
This work is partially supported by the National Natural Science Foundation of China (Nos. 62450072, U22B2052,
62302078), Central Guidance for Local Science and Technology Development Fund (Youth Science Fund Project, Category A, No. 2025JH6/101100001), the Distinguished
Young Scholars Funds of Dalian (No. 2024RJ002), the Chi-
na Postdoctoral Science Foundation (No. 2023M730741) and the Fundamental Research Funds for the Central Universities.